A uniform magnetic field of 0.20 × 10^-3 T exists in the space. Find the change in the magnetic scalar potential as one moves through 50 cm along the field.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. B = - dv/dℓ ⇒ dv = -B dℓ = - 0.2 * 10^-3 * 0.5 = - * 10^-3 T-m Since the sigh is –ve therefore potential decreases.